# Evaluating AI systems under uncertain ground truth: a case study in dermatology

### v1

@@ B.1 Modeling annotations as partial rankings
with the last block $b_{L}=b_{3}$ capturing all unranked classes. We say that annotators provide a full ranking if all blocks except $b_{L}$ have size one, i.e., $|b_{i}|=1$, $i<L$. We say the annotators provide a complete ranking if all blocks are of size one, i.e., $|b_{i}|=1$, $i\leq L$. Note that we can have multiple annotators providing partial rankings, in which case we let $b^{r}$, $r\in[R]$, denote the annotation by the annotator $r$ out of $R$ annotators.
B.2 Inverse rank normalized (IRN) aggregation
An intuitive method for aggregating differential diagnoses to arrive at a plausibility point estimate is inverse rank normalization (IRN) (Liu et al., 2020). Given a partial ranking $b$ from a single annotator with $L$ blocks as in Equation (13), we define the unnormalized IRN score of a condition $y$ as
$\text{unnormalizedIRN}(y;b)=\sum_{i=1}^{L-1}\frac{\nicefrac{{1}}{{i}}}{|b_{{}{i}{}}|}\delta[y\in b_{{}{i}{}}].$ (14)
Unnormalized IRN assigns a weight $\nicefrac{{1}}{{i}}$ to each block $b_{{}{i}{}}$, $i=1,\dots,L-1$ and a weight of 00 for all the classes in the last block $b_{{}{}{L}}$ of unranked classes. The weight is distributed equally across all classes in a block, i.e., class ${\sigma_{{i}{j}}}$ obtains weight $\nicefrac{{{1/i}}}{{|b_{{}{i}{}}|}}$. Given $R$ partial rankings $b^{1},\ldots,b^{R}$ from $R$ annotators, we define the unnormalized IRN score of a condition $y$ as
$\text{unnormalizedIRN}(y;b^{1},\ldots,b^{R})=\sum_{r=1}^{R}\text{unnormalizedIRN}(y;b^{r})$ (15)
$\text{IRN}(y;b^{1},\ldots,b^{R})=\text{unnormalizedIRN}(k;b^{1},\ldots,b^{R})/\sum_{k=1}^{K}\text{unnormalizedIRN}(k;b^{1},\ldots,b^{R}).$ (16)
Note that normalization comes after aggregation rather than normalizing per annotator $r$ first and then averaging. Overall, IRN can be viewed as an adhoc point estimate of the plausibilities $\lambda$ without an explicit generative model for $p(\lambda|b,x)$. This also implies that we cannot sample plausibilities from $p(\lambda|b,x)$. In contrast, the next section describes a probabilistic version of IRN that allows to sample plausibilities.
B.3 Probabilistic IRN (PrIRN) aggregation
To derive a probabilistic variant of IRN (PrIRN), we interpret the calculation of unnormalizedIRN as a counting method for accumulating the statistics of a histogram on discrete data. Then, the normalized $\text{IRN}(y;b^{1},\ldots,b^{R})$ corresponds to a maximum likelihood estimate of the parameters of a multinomial distribution (the sampling distribution of a histogram) given the accumulated statistics. In this spirit, we define the sampling distribution $p(\lambda|b^{1},\ldots,b^{R})$ directly as a Dirichlet distribution $\mathcal{D}(\gamma\hat{\lambda})$ where $\hat{\lambda}_{y}=\text{IRN}(y;b^{1},\ldots,b^{R})$ and $\gamma>0$ (the concentration parameter) represents our reliability parameter. Here, the $\mathcal{D}(\gamma\hat{\lambda})$ distribution is the true posterior distribution of the parameters of a categorical distribution if labels were drawn i.i.d. and the observed counts would be $\gamma\hat{\lambda}$. Then, the reliability parameter $\gamma$ can be interpreted as the number of annotators: with increasing number of annotators we have less annotation uncertainty and for $\gamma\rightarrow\infty$ the Dirichlet distribution degenerates to a point mass on $\hat{\lambda}$ – the IRN estimate.
B.4 Plackett–Luce (PL) aggregation
One conceptual problem of PrIRN is that its weighting and aggregation choices seem to be arbitrary, in the sense of not being justified by an explicit generative model for the annotation process, i.e. $p(b|\lambda)$. Another alternative is defining a generative model, and inferring the plausibilities accordingly. More specifically, this would imply defining a likelihood $p(b|\lambda)$, and estimating $\lambda$ accordingly. Having defined a likelihood, one option would be obtaining a maximum likelihood estimate of $\lambda$ under the observations $b$. Another option is to define a prior $p(\lambda)$, and then infer $\lambda$ using the posterior $p(\lambda|b)$.
As we are dealing with rankings, the Plackett–Luce (PL) (Plackett, 1975; Luce, 2012) model is a natural choice for the likelihood $p(b|\lambda)$. It models how annotators, given the true plausibilities $\lambda$, draw full rankings by sampling without replacement from all conditions. Although very appropriate for our problem, unfortunately, the posterior of the plausibilities cannot be derived analytically under the Bayesian variants of this model. However, efficient Gibbs sampling algorithms (Caron and Doucet, 2010; Francois et al., 2012) and optimization procedures (Hunter, 2003) do exist. Here, we extend the Gibbs sampler of (Caron and Doucet, 2010) to partial rankings. Importantly, our sampler will have to be correct and efficient for partial rankings, i.e., with ties and unranked conditions.
In the following, we assume a single case for brevity, with the annotator being indexed by $r\in[R]$ and assume the plausibilities $\lambda_{k}$, $k\in[K]$, to be non-negative. The inferential procedure we describe can be repeated for any other case independently. We also assume that
$p({\sigma}|\lambda)=\prod_{r=1}^{R}p({\sigma^{{r}}}|\lambda),$ (17)
meaning that the annotators’ rankings are independent, conditioned on plausibilities. Then, moving to the description of $p({\sigma^{{r}}}|\lambda)$, the classical PL model assumes that each annotator $r$ reports a full ranking of classes by sampling without replacement, proportional to the plausibilities $\lambda$, such that the first class ${\sigma^{{r}}_{{}{1}}}$ is drawn with probability
$p({\sigma^{{r}}_{{}{1}}}=k)=\frac{\lambda_{k}}{Z}$
and the next choice ${\sigma^{{r}}_{{}{3}}}$ from $\{1\dots K\}\setminus\{{\sigma^{{r}}_{{}{1}}},{\sigma^{{r}}_{{}{2}}}\}$ with probability
$p({\sigma^{{r}}_{{}{3}}}=k|{\sigma^{{r}}_{{}{1}}},{\sigma^{{r}}_{{}{2}}})=\frac{\lambda_{k}}{Z-\lambda_{{\sigma^{{r}}_{{}{1}}}}-\lambda_{{\sigma^{{r}}_{{}{2}}}}}$
$\mathcal{L}_{r}(\lambda)=\log\lambda_{{\sigma^{{r}}_{{}{1}}}}-\log Z+\log\lambda_{{\sigma^{{r}}_{{}{2}}}}-\log(Z-\lambda_{{\sigma^{{r}}_{{}{1}}}})+\dots+\log\lambda_{{\sigma^{{r}}_{{}{K}}}}-\log\left(Z-\sum_{k=1}^{K-1}\lambda_{{\sigma^{{r}}_{{}{k}}}}\right).$
The inferential goal of interest is estimating the latent plausibilities $\lambda$ based on the annotations $b$ in the form of full rankings $\sigma$. As mentioned before, this can be done according to maximum likelihood by maximizing $\mathcal{L}(\lambda)$, or in a Bayesian way by assigning a prior distribution $p(\lambda)$ and inferring the posterior distribution $p(\lambda|{\sigma})\propto p(\lambda)\prod_{r}p({\sigma^{{r}}}|\lambda)$. As the maximum likelihood solution provides a point estimate of the plausibilities, similar to IRN, here we focus on the Bayesian approach.
We can also introduce the equivalent of a reliability parameter $\gamma$ for the PL distribution by defining a new tempered PL distribution with parameters $\bar{\lambda}=\lambda^{\gamma}$ and normalization $\bar{Z}=\sum_{k}\bar{\lambda}_{k}=\sum_{k}\lambda^{\gamma}_{k}$. Unfortunately, the resulting model no longer enjoys the conjugacy properties needed for deriving a Gibbs sampling algorithm. While it is possible to use more generic sampling methods such as Hamiltonian Monte Carlo (HMC) for this tempered PL model, qualitatively we obtain the same concentration effect by simply repeating the partial rankings for each annotator such that the likelihood contribution of each annotator to the posterior is $\gamma\mathcal{L}_{r}(\lambda)$ where $\gamma$ is a small integer. This also matches the intuition of the reliability parameter $\gamma$ in PrIRN from the previous section.
The PL model as introduced above defines a distribution over full rankings. While it is straightforward to compute the probability of a given full ranking under the PL distribution, partial rankings – especially with larger block sizes – provide a computational challenge as we need to consider all the permutations compatible with the partial rankings. This would amount to a computational complexity that scales factorially in the block size. In contrast, the next section describes a procedure for computing the exact likelihood for partial rankings that scales only exponentially in the block sizes.
B.4.1 Exact likelihood computation for partial rankings
@@
For a more concise exposition, we will temporarily assume that we have a single annotator providing a partial ranking $b$, and will introduce the multiple annotator formulation at the end of Section B.4.2. We see that for the partitioning $b=(b_{{}{}{1}},b_{{}{}{2}},\dots,b_{{}{}{L}})$ the probability $p(b|\lambda)=p(b_{{}{}{1}}\succ b_{{}{}{2}}\succ\dots\succ b_{{}{}{L}})$ admits the factorization
@@
Note that above we have redefined $b$ to include all unranked classes, if any, in $b_{L}$, such that $b_{1}\cup\dots\cup b_{L}=[K]$, hence we omit the term $p(b_{L}|\lambda,b_{1:L-1})=1$.
Here, the first term is the probability of drawing a permutation from the PL distribution that starts with a sequence of classes that are all in $b_{{}{}{1}}$:
@@
Remember that $\mathcal{S}(b)$ refers to all permutations compatible with Equation (13). With a slight abuse of notation, we let $\mathcal{S}(b_{{}{}{1}})$ be the set containing all possible permutations of the first block. Concatenating all classes such that $\sigma_{1:|b_{{}{}{i}}|}$ corresponds to the first block, this is essentially the probability that the prefix $\sigma_{1:|b_{{}{}{1}}|}$ is in the first block $b_{{}{}{1}}$. Now, we define the cumulative sum of block sizes as $c_{l}=|b_{1}|+|b_{2}|+\dots+|b_{l}|$ with $c_{0}=0$ and $c_{L}=K$ and get
@@
We refer to Appendix B.4.3 for a proof and an example, but illustrate the recursion in Figure 11.
Figure 11: Hasse diagram of subsets of $\{1,2,3,4\}$. To compute the probability of a partial order of the form $\{1,2,3,4\}\succ\bar{b}$ exactly under a PL distribution with parameters $\lambda$, we associate the value $\mathcal{R}(A)$ with each node $A\subset\{1,2,3,4\},A\neq\emptyset$. After initialization of the bottom nodes, that correspond to subsets with only a single element, the algorithm proceeds by computing values of $\mathcal{R}(A)$ layer by layer using the update $\mathcal{R}_{A}\leftarrow{\sum_{a\in A}\mathcal{R}({A\setminus\{a\}}})/{(\bar{Z}+\sum_{a\in A}\lambda_{a})}$ for each node. Here, the total plausibility of the classes in $\bar{b}$ is given as $\bar{Z}=\sum_{k\in\bar{b}}\lambda_{k}$. The desired probability is $p(\{1,2,3,4\}\succ\bar{b})=\Big{(}\prod_{k\in\{1,2,3,4\}}\lambda_{k}\Big{)}\mathcal{R}({\{1,2,3,4\}})$.
B.4.2 Gibbs sampling with partial rankings
Following (Caron and Doucet, 2010), we consider the following hierarchical Bayesian model for PL with partial rankings by introducing latent arrival times $\tau$:
@@ B.4.2 Gibbs sampling with partial rankings
where $\tau=\tau_{1:K}$, $\mathcal{G}(\alpha,\beta)$ is a Gamma distribution with shape $\alpha$ and rate $\beta$ and $\mathcal{E}(\rho)$ is an exponential distribution with rate $\rho$. Here, $P_{\sigma}$ denotes the $K\times K$ permutation matrix associated with $\sigma$ that would bring $\tau$ into an ascending sorted order [^4] For exposition related to Equation (18), we assume the latent random variable ${\sigma}$ to be a complete ranking, rather than a full ranking. As defined in Section B.1, complete ranking implies that ${\sigma}$ includes no ties or unranked classes.. $Q$ denotes a $L\times K$ matrix that represents the partition structure of the partial ranking and transforms a given permutation matrix $P_{\sigma}$ to a matrix where each row represents the sets associated with each block. Finally, $B$ denotes a $L\times K$ matrix representation of the partial ranking $b$. For example, the partial ranking $b=\{1,5\}\succ\{2,3\}\succ\{4\}$ corresponds to
@@
In this model, the goal is to construct a Markov chain with the target distribution $p(\lambda|B)$. However, directly sampling form the desired posterior is difficult, thus we sample from an augmented target distribution $p(\lambda,\tau,P_{\sigma}|B)$ instead, extending (Caron and Doucet, 2010) to the case of partial rankings. Specifically, Gibbs sampling constructs such a Markov chain by iteratively sampling from full conditionals. In our model, the conditionals are
@@
B.4.3 Likelihood computation for partial rankings
@@
where $\sigma_{1:3}\in\mathcal{S}(b_{1})$, $\sigma_{4:K}\in\mathcal{S}(\bar{b}_{1})$ and *** denotes concatenation. We set $b_{1}=\{1,2,3\}$ and $\bar{b}_{1}=\{4,\dots,K\}$. We let $\bar{Z}_{1}=\lambda_{4}+\dots+\lambda_{K}$. The probability of a permutation under the PL distribution is
@@
where $\sigma_{1:c_{1}}\in\mathcal{S}(b_{1})$, $\sigma_{c_{1}+1:c_{L}}\in\mathcal{S}(\bar{b}_{1})$, *** denotes concatenation and $c_{L}=K$. Under the PL distribution, a given permutation has the following probability
@@
$\mathcal{R}_{1}(A)=\left\{\begin{array}[]{cc}1&A=\emptyset,\\ \Big{(}{\sum_{a\in A}\mathcal{R}_{1}({A\setminus\{a\}}})\Big{)}/{(\bar{Z}_{1}+\sum_{a\in A}\lambda_{a})}&A\neq\emptyset,\leavevmode\nobreak\ A\in w^{b_{1}}.\end{array}\right.$ (22)
@@
$\mathcal{R}_{l}(A)=\left\{\begin{array}[]{cc}1&A=\emptyset,\\ \Big{(}{\sum_{a\in A}\mathcal{R}_{l}({A\setminus\{a\}}})\Big{)}/{(\bar{Z}_{l}+\sum_{a\in A}\lambda_{a})}&A\neq\emptyset,\leavevmode\nobreak\ A\in w^{b_{l}}.\end{array}\right.$
@@
Note that we assume $b$ to be a partitioning of $K$ classes, i.e. $c_{L}=K$, which implies $(\prod_{k\in b_{L}}\lambda_{k})\mathcal{R}_{l}(b_{l})=1$. In some practical applications, the last block $L$ might collect a possibly large number of “unranked” classes as explained in Section B.1. Therefore in such cases we can avoid some potentially expensive computation for the last block since we can say that $p(b|\lambda)=(\prod_{k\notin b_{L}}\lambda_{k})\prod_{l=1}^{L-1}\mathcal{R}_{l}(b_{l})=(\prod_{k=1}^{c_{L}}\lambda_{k})\prod_{l=1}^{L}\mathcal{R}_{l}(b_{l})$. ∎
B.5 Average overlap with partial rankings
In this section, we extend the average overlap definition of Section 2 to partial rankings. We first define average overlap (at $L$) between two full rankings $\sigma=({\sigma_{{}{1}}},\ldots,{\sigma_{{}{K}}})$ and $\sigma^{\prime}=({\sigma_{{}{1}}}^{\prime},\ldots,{\sigma_{{}{K}}}^{\prime})$ as
@@ B.5 Average overlap with partial rankings
The average overlap metric is higher when the overlap occurs between top ranks rather than lower ones; a property that has been found to be useful when comparing differential diagnoses (Eng et al., 2019). We can also express average overlap using matrix notation:
@@
paralleling its usage in Section B.4.2. For example, the permutation $\sigma=(4,3,1,2)$ is represented by
@@
Appendix C Case study details

### v2

@@ Appendix B Plackett–Luce likelihood and Gibbs sampling
Recall the definition of differential diagnoses as partial rankings in Section 2.3. Furthermore, we say that annotators provide a full ranking if all blocks except $b_{L}$ have size one, i.e., $|b_{i}|=1$, $i<L$. We say the annotators provide a complete ranking if all blocks are of size one, i.e., $|b_{i}|=1$, $i\leq L$. Note that we can have multiple annotators providing partial rankings, in which case we let $b^{r}$, $r\in[R]$, denote the annotation by the annotator $r$ out of $R$ annotators.
B.1 Exact PL likelihood computation for partial rankings
@@
For a more concise exposition, we will temporarily assume that we have a single annotator providing a partial ranking $b$, and will introduce the multiple annotator formulation at the end of Section B.2. We see that for the partitioning $b=(b_{{}{}{1}},b_{{}{}{2}},\dots,b_{{}{}{L}})$ the probability $p(b|\lambda)=p(b_{{}{}{1}}\succ b_{{}{}{2}}\succ\dots\succ b_{{}{}{L}})$ admits the factorization
@@
Note that above we have redefined $b$ to include all unranked classes, if any, in $b_{L}$, such that $b_{1}\cup\dots\cup b_{L}=[K]$ with $[K]$ being the set of all $K$] classes. Hence we omit the term $p(b_{L}|\lambda,b_{1:L-1})=1$. Here, the first term is the probability of drawing a permutation from the PL distribution that starts with a sequence of classes that are all in $b_{{}{}{1}}$:
@@
where $\mathcal{S}(b)$ refers to all permutations compatible with $b=(b_{{}{}{1}},b_{{}{}{2}},\dots,b_{{}{}{L}})$. With a slight abuse of notation, we let $\mathcal{S}(b_{{}{}{1}})$ be the set containing all possible permutations of the first block. Concatenating all classes such that $\sigma_{1:|b_{{}{}{i}}|}$ corresponds to the first block, this is essentially the probability that the prefix $\sigma_{1:|b_{{}{}{1}}|}$ is in the first block $b_{{}{}{1}}$. Now, we define the cumulative sum of block sizes as $c_{l}=|b_{1}|+|b_{2}|+\dots+|b_{l}|$ with $c_{0}=0$ and $c_{L}=K$ and get
@@
We refer to B.3 for a proof and an example, but illustrate the recursion in Figure 12.
Fig. 12: Hasse diagram of subsets of $\{1,2,3,4\}$. To compute the probability of a partial order of the form $\{1,2,3,4\}\succ\bar{b}$ exactly under a PL distribution with parameters $\lambda$, we associate the value $\mathcal{R}(A)$ with each node $A\subset\{1,2,3,4\},A\neq\emptyset$. After initialization of the bottom nodes, that correspond to subsets with only a single element, the algorithm proceeds by computing values of $\mathcal{R}(A)$ layer by layer using the update $\mathcal{R}_{A}\leftarrow{\sum_{a\in A}\mathcal{R}({A\setminus\{a\}}})/{(\bar{Z}+\sum_{a\in A}\lambda_{a})}$ for each node. Here, the total plausibility of the classes in $\bar{b}$ is given as $\bar{Z}=\sum_{k\in\bar{b}}\lambda_{k}$. The desired probability is $p(\{1,2,3,4\}\succ\bar{b})=\Big{(}\prod_{k\in\{1,2,3,4\}}\lambda_{k}\Big{)}\mathcal{R}({\{1,2,3,4\}})$.
B.2 PL Gibbs sampling with partial rankings
Following [15], we consider the following hierarchical Bayesian model for PL with partial rankings by introducing latent arrival times $\tau$:
@@ B.2 PL Gibbs sampling with partial rankings
where $\tau=\tau_{1:K}$, $\mathcal{G}(\alpha,\beta)$ is a Gamma distribution with shape $\alpha$ and rate $\beta$ and $\mathcal{E}(\rho)$ is an exponential distribution with rate $\rho$. Here, $P_{\sigma}$ denotes the $K\times K$ permutation matrix associated with $\sigma$ that would bring $\tau$ into an ascending sorted order [^11] For exposition related to Equation (21), we assume the latent random variable ${\sigma}$ to be a complete ranking, rather than a full ranking. As defined in Section 2.3, complete ranking implies that ${\sigma}$ includes no ties or unranked classes.. $Q$ denotes a $L\times K$ matrix that represents the partition structure of the partial ranking and transforms a given permutation matrix $P_{\sigma}$ to a matrix where each row represents the sets associated with each block. Finally, $B$ denotes a $L\times K$ matrix representation of the partial ranking $b$. For example, the partial ranking $b=\{1,5\}\succ\{2,3\}\succ\{4\}$ corresponds to
@@
In this model, the goal is to construct a Markov chain with the target distribution $p(\lambda|B)$. However, directly sampling form the desired posterior is difficult, thus we sample from an augmented target distribution $p(\lambda,\tau,P_{\sigma}|B)$ instead, extending [15] to the case of partial rankings. Specifically, Gibbs sampling constructs such a Markov chain by iteratively sampling from full conditionals. In our model, the conditionals are
@@
B.3 Likelihood computation for partial rankings
@@
where $\sigma_{1:3}\in\mathcal{S}(b_{1})$, $\sigma_{4:K}\in\mathcal{S}(\bar{b}_{1})$ and * denotes concatenation. We set $b_{1}=\{1,2,3\}$ and $\bar{b}_{1}=\{4,\dots,K\}$. We let $\bar{Z}_{1}=\lambda_{4}+\dots+\lambda_{K}$. The probability of a permutation under the PL distribution is
@@
where $\sigma_{1:c_{1}}\in\mathcal{S}(b_{1})$, $\sigma_{c_{1}+1:c_{L}}\in\mathcal{S}(\bar{b}_{1})$, * denotes concatenation and $c_{L}=K$. Under the PL distribution, a given permutation has the following probability
@@
$\mathcal{R}_{1}(A)=\left\{\begin{array}[]{cc}1&A=\emptyset,\\ \Big{(}{\sum_{a\in A}\mathcal{R}_{1}({A\setminus\{a\}}})\Big{)}/{(\bar{Z}_{1}+\sum_{a\in A}\lambda_{a})}&A\neq\emptyset,~{}A\in w^{b_{1}}.\end{array}\right.$ (25)
@@
$\mathcal{R}_{l}(A)=\left\{\begin{array}[]{cc}1&A=\emptyset,\\ \Big{(}{\sum_{a\in A}\mathcal{R}_{l}({A\setminus\{a\}}})\Big{)}/{(\bar{Z}_{l}+\sum_{a\in A}\lambda_{a})}&A\neq\emptyset,~{}A\in w^{b_{l}}.\end{array}\right.$
@@
Note that we assume $b$ to be a partitioning of $K$ classes, i.e. $c_{L}=K$, which implies $(\prod_{k\in b_{L}}\lambda_{k})\mathcal{R}_{l}(b_{l})=1$. In some practical applications, the last block $L$ might collect a possibly large number of “unranked” classes as explained in Section 2.3. Therefore in such cases we can avoid some potentially expensive computation for the last block since we can say that $p(b|\lambda)=(\prod_{k\notin b_{L}}\lambda_{k})\prod_{l=1}^{L-1}\mathcal{R}_{l}(b_{l})=(\prod_{k=1}^{c_{L}}\lambda_{k})\prod_{l=1}^{L}\mathcal{R}_{l}(b_{l})$.
Appendix C Average overlap with partial rankings
In this section, we extend the average overlap definition of Section 2.5 to partial rankings. We first define average overlap (at $L$) between two full rankings $\sigma=({\sigma_{{}{1}}},\ldots,{\sigma_{{}{K}}})$ and $\sigma^{\prime}=({\sigma_{{}{1}}}^{\prime},\ldots,{\sigma_{{}{K}}}^{\prime})$ as
@@ Appendix C Average overlap with partial rankings
The average overlap metric is higher when the overlap occurs between top ranks rather than lower ones; a property that has been found to be useful when comparing differential diagnoses [30]. We can also express average overlap using matrix notation:
@@
paralleling its usage in Section B.2. For example, the permutation $\sigma=(4,3,1,2)$ is represented by
@@
Appendix D Case study details